# Liquidity Fragmentation or Optimization? Analyzing Automated Market Makers Across Ethereum and Rollups

### v2

@@ Abstract
Layer-2 (L2) blockchains offer security guarantees for Ethereum while reducing transaction (gas) fees. Consequently, they are gaining popularity among traders at Automated Market Makers (AMMs), but Liquidity Providers (LPs) are lagging behind. Our empirical results show that AMM liquidity pools on Ethereum are oversubscribed compared to their counterparties on L2s and deliver lower returns than staking ETH. LPs would receive higher rewards by reallocating over 2/3 of the liquidity to AMMs on L2s, or staking. We employ Lagrangian optimization to find the optimal liquidity allocation strategy that maximizes LP’s rewards. Moreover, we show that the returns from liquidity provisions converge to the staking rate, and in equilibrium, liquidity provisions to any AMM should provide returns equal to staking rewards. Lastly, we measure the elasticity of trading volume with respect to TVL at AMM pools and found that at the well established blockchains an increase in TVL is not associated with an increase in trading volume.
@@ 1 Introduction
Automated Market Makers (AMMs) [32], pioneered by Uniswap on Ethereum in 2018 [3], are the foundation of Decentralized Exchanges (DEXs). DEXs enable token exchanges without counterparty risk in an atomic blockchain transaction. They were originally introduced to avoid the inefficiencies of the on-chain order book associated with high on-chain storage costs and security vulnerabilities [16]. At AMM-DEX, exchange rates are set by a conservation function and current token reserves in the AMM pool. Liquidity Providers (LPs) supply tokens, earning trading fees paid by traders. As a result, a higher trading volume results in increased fees, which are subsequently distributed among LPs. Furthermore, a higher TVL in liquidity pools attracts traders by reducing price impact during swaps, a primary cost component incurred during swap transactions [2].
@@
Nevertheless, it is not always the optimal strategy for LPs to allocate their liquidity to the AMM pools with the highest trading volume. The trading fees accrued from traders are distributed among LPs in proportion to their respective contributions to the pool’s TVL. Therefore, it is the ratio of fees to TVL, often referred to as capital efficiency, that should be the primary consideration for LPs. LPs must also carefully select token pairs in the AMM pools. Volatile crypto-currency pairs lead to impermanent loss, reducing LP rewards [14]. Arbitrageurs equalizing token prices across AMM-DEXs, or with CEXs amplify LPs’ losses from market movement [20]. The Loss-Versus-Rebalancing (LVR) metric measures LP rewards relative to arbitrageurs [20]. Finally, should the rewards from liquidity provisions fall below those achieved by staking ETH, which is widely regarded the safest token allocation to yield rewards [11], the rationale behind the liquidity provision becomes questionable. The Loss-Versus-Staking metric compares LP returns to staking rewards [11].
@@
The advent of rollups, a Layer-2 (L2) scaling solution, has shifted DeFi activities from Ethereum to its rollups [26]. The success of rollups lies in their data integrity, which is secured by staked ETH, and the significant reduction in gas fees achieved by offloading computations off-chain [25, 10]. After the Ethereum Duncan upgrade in March 2024, the swap fees on L2s have dropped below \$0.01 per transaction. Consequently, the volume of swap transactions on rollups has surpassed that on the Ethereum mainnet, albeit with lower trading volumes. The reduced gas fees on rollups have facilitated transactions within the \$1-\$10 range, a previously unfeasible scenario on Ethereum. Furthermore, most rollups support the Ethereum Virtual Machine (EVM), allowing for the seamless deployment of AMM-DEXs originally designed for Ethereum onto rollups. Currently, DEXs such as Uniswap, Curve, and their forks have been deployed on rollups, enhancing the diversity of AMM pools accessible to LPs. As a result, LPs have more options to consider for providing liquidity, which serves as the starting point for this research.
@@
Related Work.  Xu et al. [32] offers a comprehensive review of major AMM categories in their SoK. The empirical study of liquidity provisions to AMMs has primarily focused on the Ethereum blockchain, especially for Uniswap (v2) [15] and Uniswap (v3) [8, 14, 19, 22, 21]. These investigations emphasize impermanent loss as a key indicator for assessing LP profitability across different cryptocurrency pairs. Furthermore, Tiruviluamala et al. [28] proposed a framework to address impermanent loss, a major risk for LPs.
@@
1.0.1 Contribution.
This research analyzes both empirically and theoretically the liquidity provision for AMMs on L2s. We assess LP rewards on Uniswap (v3) within Ethereum and its optimistic rollups (Arbitrum, Base, Optimism), and ZK rollups (ZKsync). The contributions of this work are as follows:
  • Using Lagrangian optimization, we find the optimal allocation of liquidity across staking and AMMs pools on Ethereum and its rollups, with the objective of maximizing LP rewards.
@@ 2.1 Layer-2 Blockchain Scaling
According to the blockchain scalability trilemma [23], blockchains can prioritize only two of three: decentralization, security, or scalability. Ethereum, the main platform for DeFi with the highest Total Value Locked (TVL), prioritizes decentralization and security. This led to network congestion, high gas fees, and throughput limited to 12 transactions per second (TPS). The Layer 1 (L1) and Layer 2 (L2) scaling solutions address blockchain scalability. L1 scaling involves the development of new blockchains with novel consensus mechanisms [18], sharding [30], and their own physical infrastructure. In contrast, in L2 scaling intensive computations are executed off-chain, with their results being recorded on the underlying L1 blockchain [25, 10].
Rollup.
Rollups [27], the most popular form of L2 scaling, act as blockchains. They generate blocks, execute transactions, and subsequently record compressed batched data on the L1. This approach ensures that the integrity of the rollup data is guaranteed by L1 security, such as staked ETH in the case of Ethereum’s rollups. In order to maliciously modify the rollup state history, an attacker must compromise the security of the underlying L1 network.
Sequencer.
A sequencer [5] is an integral component of rollup, tasked with executing and ordering transactions, forming blocks, and creating batches that are uploaded to the L1 chain. By bundling transactions together, rollups provide more gas-efficiency compared to the L1 network. To avoid additional trust assumptions in sequencer operators, rollups use optimistic or zero-knowledge proofs (ZKPs) to ensure the correctness and integrity of the data. Fig. 1 illustrates the architecture of optymistic and ZK-rollups. Presently, major Ethereum optimistic rollups (e.g., Arbitrum, Optimism, Base) and ZK-rollups (e.g., ZKsync, StarkNet) rely on centralized sequencers and, in the case of ZK-rollup, centralized provers. Ongoing research aims to decentralize sequencers [24] and provers [31] to improve the reliability and security of both types of rollups.
Optimistic Rollup.
Optimistic rollups [17] operate on a trust-based system, assuming transactions are valid unless disputed. This approach simplifies the implementation of optimistic rollups, particularly in supporting the Ethereum Virtual Machine (EVM). Consequently, optimistic roll-ups were faster to launch EVM compatibility and attracted higher DeFi adoption. However, the optimistic fraud-proof mechanism can lead to delays in withdrawals. Currently, most optimistic rollups enforce a 7-day challenge period.
ZK-Rollup.
In contrast, ZK-rollups [6] leverage ZKPs to validate the state on L1 immediately after a proof has been generated off-chain by provers and submitted to verifiers. Verifiers, smart contracts on L1, validate transactions aggregated by the sequencer and confirm their correctness. Consequently, rapid finality is ensured, albeit for increased computational costs. ZK-rollups also offer enhanced compression opportunities, e.g. by posting to L1 only ZKPs instead of all transaction data.
@@ 2.2 Automated Market Maker
Constant Function Market Makers (CFMM) represent the main type of AMM and are characterized by a trading functioncommonly referred to as a conservation function or reserve curve. This trading function maps token reserves to the invariant, which remains constant throughout each swap transaction. The number of tokens within the pool, $N$, is typically $N=2$. The token reserves within liquidity pools are represented by $x_{1},\ldots,x_{N}$. A diverse array of CFMMs has been proposed [32] to mitigate the trading cost (price impact) for traders and maximize capital efficiency for LPs.
@@ 3 Model
In our model, we also consider that LPs have the possibility to stake and earn the staking rewards $r$. We consider the staking rate $r_{s}$ to be similar to the risk-free rate in traditional finance. Consequently, we assume that if liquidity pools do not provide higher rewards than the staking $r_{s}$, the LP reallocates her wealth to staking. Given that, within the Ethereum ecosystem, the volume of capital allocated to staking vastly exceeds that which is locked within liquidity pools, we assume that the staking rate remains unaffected by staked capital in our model.
@@ 3.1 Constant Product Market Maker
$fees_{i}\frac{w_{i}}{\text{TVL}_{i}+w_{i}}=f\cdot\text{Vol}_{i}\frac{w_{i}}{\text{TVL}_{i}+w_{i}}$ (1)
Thus, the return rate of LP for allocating $w_{i}$ capital to the liquidity pool on blockchain $i$ is:
@@
Returns, as presented by DEX aggregators and the GUI of AMM-DEX, are $r_{i}(0)$.
@@ 3.3 Optimal Allocation
Assume there is one LP. We denote by $w_{0}$ the amount of capital allocated to staking at the staking rate $r_{s}$ and for $i=1,\ldots,n$ we denote $w_{i}$ as the capital allocated to liquidity pool $i$. We are looking for a vector $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n})$ so that all $r_{i}(w_{i})$ are equal for $i>0$. We assume that the LP has total liquidity $W$ to allocate to staking and AMM pools. By allocating $w_{0}$ to staking, the return from staking is $r_{s}w_{0}$.
@@
$\max_{\mathbf{w}=(w_{0},\ldots,w_{n})}r_{s}w_{0}+\sum_{i=1}^{n}\frac{f\cdot\text{Vol}_{i}}{\text{TVL}_{i}+w_{i}}w_{i}$ (4)
@@ 3.4 Convergence
Assume now that there are $m$ LPs that invest into the staking and AMMs consecutively. An allocation of capital to a pool increases its TVL and thereby decreases the rate of return, see Equation (2). If LPs allocate their capital consecutively, the rate of return of each pool will decrease until it reaches $r_{s}$. Once each pool reaches $r_{s}$, the remaining LPs will invest into staking. Formally, we have the following convergence result. Let $r^{j}_{i}$ be the rate of return for pool $i$, after $j-1$ LPs have invested their capital. Note that after each investment by an LP, the respective pool’s TVLs increase by that amount.
@@ 3.5 Modeling Volume Dynamics
We introduce a parameter for volume elasticity, denoted by $\epsilon_{v}$, to capture the relationship between trading volume (Vol) and TVL in the pool:
@@
where $k$ is a scaling constant, and $\epsilon_{v}$ represents the elasticity of trading volume with respect to TVL.
@@
$\max_{\mathbf{w}=(w_{0},\ldots,w_{n})}r_{s}w_{0}+\sum_{i=1}^{n}\frac{f\cdot k_{i}\cdot(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-1}}{\text{TVL}_{i}+w_{i}}w_{i},$ (7)
subject to $w_{i}\geq 0$ for each $i=0,\ldots,n$ and $\sum_{i=0}^{n}w_{i}=W$.
@@
The allocation vector $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n})$ with
$w_{i}=\text{TVL}_{i}(\sqrt{f\cdot k_{i}\cdot(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-1}r_{s}^{-1}}-1),\quad i=1,\ldots,n,$ (8)
and $w_{0}=W-\sum_{i=1}^{n}w_{i},$ is the solution to the optimization problem (7) and maximizes LP’s earnings.
@@
The statement is proved as in Proposition 1 using the Lagrangian $\mathcal{L}(\mathbf{w},\lambda)=r_{s}w_{0}+\sum_{i=1}^{n}\frac{f\cdot k_{i}\cdot(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-1}}{\text{TVL}_{i}+w_{i}}w_{i}+\lambda\left(W-\sum_{i=0}^{n}w_{i}\right).$
@@
Assume $m$ LPs invest sequentially into staking and AMM pools. An allocation of capital to a pool increases its TVL and reduces the rate of return. The rate of return for pool $i$ after $j-1$ LPs have invested is given by:
@@
The return from each pool converges to the staking rate $r_{s}$ as $m\to\infty$. That is, $r_{i}^{m}(w_{i}^{m})\to r_{s}\quad\text{as }m\to\infty\quad\text{for all }i=1,\ldots,n.$

### v3

@@ Abstract
Layer-2 (L2) blockchains inherit Ethereum’s security guarantees while reducing gas fees. As a result, they are gaining traction among traders at Automated Market Makers (AMMs), sparking debate over whether they contribute to liquidity fragmentation of Ethereum. Our research suggests that such fragmentation is not currently occurring. However, it could emerge in the future—particularly if Liquidity Providers (LPs) recognize the higher returns available on L2s. Using Lagrangian optimization, we develop a model for optimal liquidity allocation across AMMs on Ethereum and its L2s, using staking as a benchmark. We show that, in equilibrium, AMM liquidity provision returns converge to this reference rate. Additionally, we measure the elasticity of trading volume with respect to Total Value Locked (TVL) in AMMs and find that, on well-established blockchains, an increase in TVL does not necessarily lead to higher trading volume. Finally, our empirical findings reveal that Ethereum’s liquidity pools are oversubscribed compared to those on L2s and often yield lower returns than staking Ether. LPs could maximize their rewards by reallocating more than two-thirds of their liquidity to L2s and staking.
@@ 1 Introduction
Automated Market Makers (AMMs) [37], pioneered by Uniswap on Ethereum in 2018 [4], are the foundation of Decentralized Exchanges (DEXs). DEXs enable token exchanges without counterparty risk in an atomic blockchain transaction [15]. They were originally introduced to avoid the inefficiencies of the on-chain order book associated with high on-chain storage costs and security vulnerabilities [22]. At AMM-DEX, exchange rates are set by a conservation function and current token reserves in the AMM pool. Liquidity Providers (LPs) supply tokens, earning trading fees paid by traders. As a result, a higher trading volume results in increased fees, which are subsequently distributed among LPs. Furthermore, a higher Total Value Locked (TVL) in liquidity pools attracts traders by reducing price impact during swaps, a primary cost component incurred during swap transactions [3].
@@
Nevertheless, it is not always the optimal strategy for LPs to allocate their liquidity to the AMM pools with the highest trading volume. The trading fees accrued from traders are distributed among LPs in proportion to their respective contributions to the pool’s TVL. Therefore, it is the ratio of fees to TVL, often referred to as capital efficiency, that should be the primary consideration for LPs. LPs must also carefully select token pairs in the AMM pools. Volatile crypto-currency pairs lead to impermanent loss, reducing LP rewards [19]. Arbitrageurs equalizing token prices across AMM-DEXs, or with Centralized Exchanges (CEXs) amplify LPs’ losses from market movement [26]. The Loss-Versus-Rebalancing (LVR) metric measures LP rewards relative to arbitrageurs [26]. Finally, should the rewards from liquidity provisions fall below those achieved by staking ETH, which is widely regarded the safest token allocation to yield rewards [14], the rationale behind the liquidity provision becomes questionable. The Loss-Versus-Staking metric compares LP returns to staking rewards [14].
@@
The advent of rollups, a Layer-2 (L2) scaling solution [32], has shifted DeFi activities from Ethereum to its rollups. The success of rollups lies in their data integrity, which is secured by staked ETH, and the significant reduction in gas fees achieved by offloading computations off-chain [31, 13]. After the Ethereum Dencun upgrade in March 2024, the swap fees on L2s have dropped below \$0.01 per transaction. Consequently, the volume of swap transactions on rollups has surpassed that on the Ethereum mainnet, albeit with lower trading volumes. The reduced gas fees on rollups have facilitated transactions within the \$1-\$10 range, a previously unfeasible scenario on Ethereum. Furthermore, most rollups support the Ethereum Virtual Machine (EVM), allowing for the seamless deployment of AMM-DEXs originally designed for Ethereum onto rollups. Currently, DEXs such as Uniswap, Curve, and their forks have been deployed on rollups, enhancing the diversity of AMM pools accessible to LPs. As a result, LPs have more options to consider for providing liquidity, which serves as the starting point for this research.
@@
Related Work.  A comprehensive review of major AMM categories has been made available in [37] and expanded and updated in [30]. The empirical study of liquidity provisions to AMMs has primarily focused on the Ethereum blockchain, especially for Uniswap (v2) [20] and Uniswap (v3) [11, 19, 25, 28, 27]. These investigations emphasize impermanent loss as a key indicator for assessing LP profitability across different cryptocurrency pairs. Furthermore, Tiruviluamala et al. [34] proposed a framework to address impermanent loss.
@@
Contribution.  This research analyzes both empirically and theoretically the liquidity provision for AMMs on L2s. We assess LP rewards on Uniswap (v3) within Ethereum and its optimistic rollups (Arbitrum, Base, Optimism), and ZK rollups (ZKsync). The contributions of this work are as follows:
  • Using Lagrangian optimization, we find the optimal allocation of liquidity across staking and AMMs pools on Ethereum and its rollups, with the objective of maximizing LP rewards. We further show that in equilibrium, liquidity provisions to AMMs should provide returns equal to risk-free rate across all blockchains.
@@ 2.1 Layer-2 Blockchain Scaling
According to the blockchain scalability trilemma [29], blockchains can prioritize only two of three: decentralization, security, or scalability. Ethereum, the main platform for DeFi with the highest TVL, prioritizes decentralization and security. This led to network congestion, high gas fees, and throughput limited to 12 transactions per second (TPS). The Layer 1 (L1) and Layer 2 (L2) scaling solutions address blockchain scalability. L1 scaling involves the development of new blockchains with novel consensus mechanisms [24], sharding [36], and their own physical infrastructure. In contrast, in L2 scaling intensive computations are executed off-chain, with their results being recorded on the underlying L1 blockchain [31, 13].
Rollup.  Rollups [33], the non-custodial form of L2 scaling, act as blockchains. They generate blocks, execute transactions, and subsequently record compressed batched data on the L1. This approach ensures that the integrity of the rollup data is guaranteed by L1 security, such as staked ETH in the case of Ethereum’s rollups. In order to maliciously modify the rollup state history, an attacker must compromise the security of the underlying L1 network.
Sequencer.  A sequencer [6] is an integral component of rollup, tasked with executing and ordering transactions, forming blocks, and creating batches that are uploaded to the L1 chain. By bundling transactions together, rollups provide more gas-efficiency compared to the L1 network. To avoid additional trust assumptions in sequencer operators, rollups use optimistic or zero-knowledge proofs (ZKPs) to ensure the correctness and integrity of the data. Fig. 1 illustrates the architecture of optymistic and ZK-rollups. Presently, major Ethereum optimistic rollups (e.g., Arbitrum, Optimism, Base) and ZK-rollups (e.g., ZKsync, StarkNet) rely on centralized sequencers and, in the case of ZK-rollup, centralized provers.
Optimistic Rollup.  Optimistic rollups [23] operate on a trust-based system, assuming transactions are valid unless disputed. This approach simplifies the implementation of optimistic rollups, particularly in supporting the Ethereum Virtual Machine (EVM). Consequently, optimistic roll-ups were faster to launch EVM compatibility and attracted higher DeFi adoption. However, the optimistic fraud-proof mechanism can lead to delays in withdrawals. Currently, most optimistic rollups enforce a 7-day challenge period.
ZK-Rollup.  In contrast, ZK-rollups [7] leverage ZKPs to validate the state on L1 immediately after a proof has been generated off-chain by provers and submitted to verifiers. Verifiers, smart contracts on L1, validate transactions aggregated by the sequencer and confirm their correctness. Consequently, rapid finality is ensured, albeit for increased computational costs. ZK-rollups also offer enhanced compression opportunities, e.g. by posting to L1 only ZKPs instead of all transaction data.
@@ 2.2 Automated Market Maker
Constant Function Market Makers (CFMM) represent the main type of AMM and are characterized by a trading function, commonly referred to as a conservation function or reserve curve. This trading function maps token reserves to the invariant, which remains constant throughout each swap transaction. The number of tokens within the pool, $N$, is typically $N=2$. The token reserves within liquidity pools are represented by $x_{1},\ldots,x_{N}$. A diverse array of CFMMs has been proposed [37] to mitigate the trading cost (price impact) for traders and maximize capital efficiency for LPs.
@@ 3 Model
  • Each liquidity pool $i$ consists of the same tokens and consequently has the same impermanent loss (IL).
@@
In our model, we also consider that LPs have the possibility to stake and earn the staking rewards $r$. We consider the staking rate $r_{s}$ to be similar to the risk-free rate in traditional finance. Consequently, we assume that if liquidity pools do not provide higher rewards than the staking $r_{s}$, the LP reallocates her wealth to staking. Given that, within the Ethereum ecosystem, the volume of capital allocated to staking vastly exceeds that which is locked within liquidity pools [18], we assume that the staking rate remains unaffected by staked capital in our model.
@@ 3.1 Constant Product Market Maker
$\frac{w_{i}}{\text{TVL}_{i}+w_{i}}fees_{i}=\frac{w_{i}}{\text{TVL}_{i}+w_{i}}f\cdot\text{Vol}_{i}$ (1)
Thus, the return rate of LP on fees for allocating $w_{i}$ capital to the liquidity pool on blockchain $i$ is:
@@
Returns, as presented by DEX aggregators and the GUI of AMM-DEX, are $r_{i}(0)$. In order to calculate the LP return, impermanent loss (IL) must be deduced from $r_{i}(0)$.
@@ 3.3 Optimal Allocation
Assume there is one LP. We denote by $w_{0}$ the amount of capital allocated to staking at the staking rate $r_{s}$ and for $i=1,\ldots,n$ we denote $w_{i}$ as the capital allocated to liquidity pool $i$. We are looking for a vector $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n})$ that maximizes LP’s earnings. We assume that the LP has total liquidity $W$ to allocate to staking and AMM pools. By allocating $w_{0}$ to staking, the return from staking is $r_{s}w_{0}$.
@@
$\max_{\mathbf{w}=(w_{0},\ldots,w_{n})}r_{s}w_{0}+\sum_{i=1}^{n}\frac{w_{i}}{\text{TVL}_{i}+w_{i}}f\cdot\text{Vol}_{i}$ (4)
@@
Incorporating impermanent loss into the above model can easily be done by increasing the reference rate $r_{s}$ with the rate of impermanent loss, as each liquidity pool incurs an equal impermanent loss.
@@ 3.4 Convergence
Assume now that there are $m$ independent and identical LPs that invest into the staking and AMMs consecutively. Each action of an LP changes the return rates of the AMMs. Based on these changes other LPs will take action accordingly. In particular, an allocation of capital to a pool increases its TVL and thereby decreases the rate of return, see Equation (2). If LPs allocate their capital consecutively, the rate of return of each pool will decrease until it reaches $r_{s}$. Once each pool reaches $r_{s}$, the remaining LPs will invest into staking. Formally, we have the following convergence result. Let $r^{j}_{i}$ be the rate of return for pool $i$, after $j-1$ LPs have invested their capital. Note that after each investment by an LP, the respective pool’s TVLs increase by that amount.
@@ 3.5 Modeling Volume Dynamics
We introduce a parameter for volume elasticity [21], denoted by $\epsilon_{v}$. The elasticity parameter captures how sensitive trading activity is to changes in liquidity. The trading volume follows the functional form
@@
where $k$ is a scaling constant. This form is particularly motivated because deeper liquidity pools (i.e. pools with higher liquidity) reduce price impact and slippage, incentivizing larger trades.
@@
$\max_{\mathbf{w}=(w_{0},\ldots,w_{n})}r_{s}w_{0}+\sum_{i=1}^{n}f\cdot k_{i}\cdot(\text{TVL}_{i}+w_{i})^{\epsilon_{v}-1}w_{i}$ (7)
subject to $w_{i}\geq 0$ for each $i=0,\ldots,n$ and $\sum_{i=0}^{n}w_{i}=W$. Note that we replaced $\text{Vol}_{i}$ with (6), considering that the LP’s contribution $w_{i}$ to pool $i$ must be added to $\text{TVL}_{i}$.
@@
The allocation vector $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n})$ which solves
$w_{i}=\left(\frac{r_{s}}{fk_{i}(\text{TVL}_{i}+\epsilon_{v}w_{i})}\right)^{\frac{1}{\epsilon_{v}-2}}-\text{TVL}_{i},\quad i=1,\ldots,n,$ (8)
and $w_{0}=W-\sum_{i=1}^{n}w_{i},$ is the solution to the optimization problem (7) and maximizes LP’s earnings. For $w_{i}$ small with respect to TVLi, the above equation can be approximated:
$w_{i}\approx\left(\frac{r_{s}}{fk_{i}\text{TVL}_{i}}\right)^{\frac{1}{\epsilon_{v}-2}}-\text{TVL}_{i},\quad i=1,\ldots,n.$ (9)
@@
The proof is included in Appendix 0.A.
The optimal allocation can be numerically calculated using (8). An approximate result for small $w_{i}$ is given in (9).
@@
Assume $m$ identical independent LPs invest sequentially into staking and AMM pools. An allocation of capital to a pool increases its TVL and reduces the rate of return. The rate of return for pool $i$ after $j-1$ LPs have invested is given by:
@@
For $\epsilon_{v}<1$, the return from each pool converges to the staking rate $r_{s}$ as $m\to\infty$. That is, $r_{i}^{m}(w_{i}^{m})\to r_{s}\quad\text{as }m\to\infty\quad\text{for all }i=1,\ldots,n.$
@@
Note that if $\epsilon_{v}>1$, increasing TVL increases the return instead of reducing it. As more LPs enter, trading volume scales more than liquidity, leading to increasing LP rewards rather than convergence to $r_{s}$. This creates a self-reinforcing effect, where AMM pools can keep attracting liquidity as returns remain high, contradicting the equilibrium assumption. This implies that LPs should always prefer AMM liquidity over staking when $\epsilon_{v}>1$, as staking will always yield lower returns. For $\epsilon_{v}=1$, the rates of return are constant for each pool.